OpenStudy (anonymous):

Which value of x makes the equation true? x – 5 = –13 A. 18 B. 8 C. –8 D. –18

OpenStudy (anonymous):

x-5=-13 -> we need to isolate x x-5+5 = -13+5 ->Add 5 to both sides to cancel out the -5 x = -8

OpenStudy (ksaimouli):

-8

OpenStudy (anonymous):

x-5=-13 x-5+5=-13+5 x=-8 So, yhe answer is C